OpenStudy (silverlark2):

Use the substitution y = u- 2x to find the general solution of the equation dy/dx = -(8x + 4y +1) / (4x + 2y+ 1)

OpenStudy (priyar):

did u substitute and try to solve?

OpenStudy (priyar):

what do u get after substituting y=u-2x in the expression?

OpenStudy (priyar):

@Silverlark2 ?

OpenStudy (baru):

u=y+2x so 8x+4y=4u and 4x+2y= 2u

OpenStudy (baru):

differentiate y=u-2x wrt x\[\frac{dy}{dx}=\frac{du}{dx}-2\]

OpenStudy (priyar):

ya this is a simpler way..

OpenStudy (baru):

make all these substitutions and integrate

OpenStudy (silverlark2):

ok i think i got it! thanks to both :)

OpenStudy (priyar):

welcome!